OpenStudy (anonymous):

Which value of x makes the following equation true? 5.6(x - 3) = -0.4x + 19.2 9 6 5 8

OpenStudy (the_fizicx99):

5.6x - 16.8 = -.4x + 19.2 +16.8 +16.8 5.6x = -.4x + 36 +.4x +.4x 6x = 36 /6 /6 x = 6

OpenStudy (anonymous):

THANKKKK YOUUUU!!;DDD ;)

OpenStudy (the_fizicx99):

Just distribute and solve, combine like terms. :p